OpenStudy (anonymous):

3.5b-33.4=75.1 need a solution

OpenStudy (katrinakaif):

3.5b = 108.5 Divide by 3.5 = b = 31

jimthompson5910 (jim_thompson5910):

3.5b-33.4=75.1 35b-334=751 ... Multiply EVERY term by 10 to make every number a whole number 35b=751+334 35b=1085 b = 1085/35 b = 31 So the answer is b = 31